Mickey, a daredevil mouse of mass 0.0217 kg, is attempting to become the world's first "mouse cannonball." He is loaded into a spring‑powered gun pointing up at some angle and is shot into the air. The gun's spring has a force constant of 58.9 N/m and is initially compressed a distance of 0.149 m from its relaxed position. If Mickey has a constant horizontal speed of 2.01 m/s while he is flying through the air, how high ℎ above his initial location in the gun does Mickey soar? Assume ?=9.81 m/s2. ℎ= m
Answer #1

Potential energy, P = 0.5 x 58.9 x 0.149^2 = 0.65 J

Kinetic energy, K = 0.5 x 0.0217 x 2.01^2 = 0.044 J

Height, h = (0.65 - 0.044)/(0.0217 x 9.81) = 2.85 m

Comment in case any doubt please rate my answer....